OpenStudy (van.gosh):

Express the complex number in trigonometric form: 1-i

OpenStudy (anonymous):

let \[1-\iota =r \left( \cos \theta +\iota \sin \theta \right)\] \[r \cos \theta=1,r \sin \theta=-1 \] square and add \[r^2=1+1=2\] \[r=\sqrt{2}\] \[\cos \theta=\frac{ 1 }{ \sqrt{2} },\sin \theta=\frac{ -1 }{ \sqrt{2} }\] so\[\ theta\] lies in 4th quadrant.

OpenStudy (anonymous):

\[\cos \theta=\cos \left( 2n \pi+2\pi-\frac{ \pi }{ 4 } \right),\theta=?\] where n is an integer.

OpenStudy (van.gosh):

Thank you so much! I appreciate all the work it must have taken to type that out :-)

OpenStudy (van.gosh):

Is it alright if I ask a question? I'm lost on one of the steps.

OpenStudy (van.gosh):

I'm confused about how you went from the \[rcos \Theta =1\]\[rsin \Theta= -1\] to the part where you square and do \[r^{2}=1+1=2\] \[r=\sqrt{2}\]